# Computation of Greeks under rough Volterra stochastic volatility models using the Malliavin calculus approach

### v1

@@ 3.2 Rough Volterra SV models
which is an integral form of (18).
@@
$V_{t}=\varphi(t,Y_{t}),$ (30)
where $\varphi$ is a function differentiable in the second variable, or $V$ can be given as a solution to the stochastic differential equation
@@
    $\rho$ is the correlation between processes $V$ and $S$,
  3. 3.
    under the volatility or variance model (31), using Proposition 2.3 and the fact that the stochastic differential equation satisfied by $D_{s}V_{t}$ is linear, we have
    $D_{s}V_{t}=\rho v(V_{s})\mathbf{1}_{[0,t]}(s)\exp\left\{\int_{s}^{t}[u^{\prime}(V_{u})-\frac{1}{2}(v^{\prime}(V_{u}))^{2}]\,\mathrm{d}s+\int_{s}^{t}v^{\prime}(V_{u})\,\mathrm{d}Z_{u}\right\}.$ (32)
    This allow us to compute $G(t,T)$ and $D_{s}G(t,T)$ with $s\leq t,$ under different stochastic volatility models.
@@
    In the $\alpha$RFSV model, first introduced by Merino, Pospíšil, Sobotka, Sottinen, and Vives (2021), $\sigma(x)={x}$ and the process $V_{t}$ is given explicitly by (30) with
    $\varphi(t,x)=V_{0}\cdot\exp\left\{\xi\cdot x-\frac{1}{2}\alpha\xi^{2}r(t)\right\}.$
@@
    a new mixed $\alpha$RFSV model, a two factor rough model, where two fractional volatility factors are considered, one with $H<\frac{1}{2}$ and another with $H^{\prime}>\frac{1}{2}$. This model is a generalisation of the $\alpha$RFSV model (Merino, Pospíšil, Sobotka, Sottinen, and Vives, 2021) and the mixed fractional Bergomi model (Alòs and León, 2021).
@@
Let $G(t,T)$ be defined as in (20). Then the formulas for the Greek Vega and for the derivative with respect the Hurst parameter $H$ are
$\mathcal{V}$ $=\mathbb{E}\left[f(S_{T})\,\delta\left(\frac{\mathrm{e}^{-rT}\partial_{V_{0}}S_{T}}{\int_{0}^{T}D_{u}S_{T}\,\mathrm{d}u}\right)\right]$
$=\mathrm{e}^{-rT}\mathbb{E}\left[f(S_{T}^{\theta})\cdot\delta(c_{V_{0}}(V_{s}))\right]$
$\mathcal{H}$ $=\mathbb{E}\left[f(S_{T})\,\delta\left(\frac{\mathrm{e}^{-rT}\partial_{H}S_{T}}{\int_{0}^{T}D_{u}S_{T}\,\mathrm{d}u}\right)\right]$
$=\mathrm{e}^{-rT}\mathbb{E}\left[f(S_{T}^{\theta})\cdot\delta(c_{H}(V_{s}))\right],$
where $c_{V_{0}}$ and $c_{H}$ are given by (23) with $a$ and $b$ being the corresponding partial derivatives (21).
Proof.
Straightforward application of (22). ∎
4 Greeks formulas in rough Volterra SV models
In this section we calculate Greeks formulas in the first class of the Volterra SV models, in particular for the $\alpha$RFSV model (Section 4.1), for the mixed $\alpha$RFSV model (Section 4.2) and for the Stein and Stein model (Section 4.3).
@@ 4.1 $\alpha$RFSV model
To calculate the considered Greeks, we need to calculate $D_{s}V_{r}$, $D_{t}D_{s}V_{r}$, $G(t,T)$, and $D_{s}G(t,T).$
@@
$D_{t}D_{s}V_{r}=\rho^{2}\xi^{2}K_{H}(r,s)K_{H}(r,t)V_{r}\mathbf{1}_{[0,r]}(s\vee t)$
@@
Using this formulas we have
@@
$=\int_{0}^{T}V_{t}\,\mathrm{d}t+\rho\xi\left(\int_{0}^{T}V_{r}\kappa_{r}\,\mathrm{d}W_{r}-\int_{0}^{T}V_{r}^{2}\kappa_{r}\,\mathrm{d}r\right)$ (34)
and
@@
$=2\rho\xi\int_{0}^{T}V_{t}\kappa_{t}\,\mathrm{d}t+\rho^{2}\xi^{2}\left(\int_{0}^{T}V_{r}\kappa_{r}^{2}\,\mathrm{d}W_{r}-2\int_{0}^{T}V_{r}^{2}\kappa_{r}^{2}\,\mathrm{d}r\right).$ (35)
These are all necessary ingredients to plug in to the Greeks formulas in Theorem 3.3.
@@
In Figure 1 we can see a convergence of the Delta for the $\alpha$RFSV model with $H=0.14$, $\alpha=1$ (rBergomi), $V_{0}=0.62$, $\xi=0.21$ and $\rho=-0.05$. Numerically, the kernel (17) is considered with $\varepsilon=10^{-6}$. Market values are $S_{0}=100$ (dollars) and $r=0.05$, option parameters are strike price $K=100$ (dollars) and maturity $T=1$ (year). Monte-Carlo (MC) simulation point estimate together with the 99% confidence interval (CI) is plotted as a dependence on the number of simulations (NS). In particular $\Delta_{\text{call}}=0.0272\pm 0.0110$ and $\Delta_{\text{put}}=-0.0008\pm 0.0052$. Although we have no analytical $\alpha$RFSV formula for Delta, we plot the horizontal dashed line which is the numerical mean at the end (for the largest number of simulations).
Figure 1: Convergence of the $\alpha$RFSV Delta in the case $H=0.14$.

### v2

@@ 3.2 Rough Volterra SV models
which is the integral form of (24).
@@
$V_{t}=f(t,Y_{t}),$ (35)
where $f$ is a function differentiable in the second variable, or $V$ can be given as a solution to the stochastic differential equation
@@
    $\rho$ is the correlation between processes $V$ and $S$.
@@
    In the $\alpha$RFSV model, first introduced by Merino, Pospíšil, Sobotka, Sottinen, and Vives (2021), $\sigma(x)={x}$ and the variance process $V_{t}$ is given explicitly by (35) with
    $f(t,x)=V_{0}\cdot\exp\left\{\xi\cdot x-\frac{1}{2}\alpha\xi^{2}r(t)\right\}.$
@@
    In a new mixed $\alpha$RFSV model, a two factor rough model, two fractional volatility factors are considered, one with $H<\frac{1}{2}$ and another with $H^{\prime}>\frac{1}{2}$. This model is a generalisation of the $\alpha$RFSV model, see Merino, Pospíšil, Sobotka, Sottinen, and Vives (2021) and the mixed fractional Bergomi model, see Alòs and León (2021).
@@
4 Greeks formulas in rough Volterra SV models
In this section we calculate Greeks formulas for the $\alpha$RFSV model (Section 4.1), for the mixed $\alpha$RFSV model (Section 4.2) and for the Stein and Stein model (Section 4.3).
Notice that the formulas given in Theorem 3.3 depend on the process $h$ and the Skorohod integral $\delta$. In the proof of Theorem 3.3, the parameter $\varepsilon>0$ did not play an important role for the proof, but will play instead an important role in the numerics. Indeed, since we can choose $h$ such that $||\mathbf{1}_{[0,T]}(t)-h_{t}||_{L^{2}(\Omega\times[0,T])}\leq\varepsilon$ for $\varepsilon>0$ arbitrarily small, we will simulate the formulas of Theorem 3.3 for $h_{t}=\mathbf{1}_{[0,T]}(t)$. Using the Theorem 2.5, we can derive approximate formulas for the Greeks.
@@ 4 Greeks formulas in rough Volterra SV models
Let $\mathcal{G}_{T}:=\int_{0}^{T}G(u,T)\,\mathrm{d}u$. The approximate formulas for the Greeks are:
$\Delta=$ $\frac{\mathrm{e}^{-rT}}{S_{0}}\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\left(\frac{W_{T}}{\mathcal{G}_{T}}+\frac{\int_{0}^{T}D_{s}\mathcal{G}_{T}\,\mathrm{d}s}{\mathcal{G}_{T}^{2}}\right)\right],$
$\Gamma=$ $-\frac{\mathrm{e}^{-rT}}{S_{0}^{2}}\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\left(\frac{W_{T}}{\mathcal{G}_{T}}+\frac{\int_{0}^{T}D_{s}\mathcal{G}_{T}\,\mathrm{d}s}{\mathcal{G}_{T}^{2}}\right)\right]$
$+\frac{\mathrm{e}^{-rT}}{S_{0}^{2}}\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\left(\frac{W_{T}}{\mathcal{G}_{T}}+\frac{\int_{0}^{T}D_{s}\mathcal{G}_{T}\,\mathrm{d}s}{\mathcal{G}_{T}^{2}}\right)^{2}\right]$
$-\frac{\mathrm{e}^{-rT}}{S_{0}^{2}}\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\left(\frac{T}{\mathcal{G}_{T}}-\frac{W_{T}\int_{0}^{T}D_{t}\mathcal{G}_{T}\,\mathrm{d}t}{\mathcal{G}_{T}^{2}}+\frac{\int_{0}^{T}\int_{0}^{T}D_{t}D_{s}\mathcal{G}_{T}\,\mathrm{d}s\,\mathrm{d}t}{\mathcal{G}_{T}}-\frac{2\left(\int_{0}^{T}D_{s}\mathcal{G}_{T}\,\mathrm{d}s\right)^{2}}{\mathcal{G}_{T}^{3}}\right)\right],$
ϱ=italic-$\varrho=$ $rT\mathrm{e}^{-rT}\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\left(\frac{W_{t}}{\mathcal{G}_{T}}+\frac{\int_{0}^{T}D_{s}\mathcal{G}_{T}\,\mathrm{d}s}{\mathcal{G}_{T}^{2}}\right)\right]-\mathrm{e}^{-rT}\mathbb{E}_{\mathbb{Q}}[f(S_{T})],$
$\mathcal{V}=$ $\mathrm{e}^{-rT}\mathbb{E}_{\mathbb{Q}}\left[f(S_{t})\left(\frac{W_{T}\int_{0}^{T}b_{V_{0}}(V_{s})\,\mathrm{d}s}{\mathcal{G}_{T}}-\frac{\int_{0}^{T}\int_{0}^{T}D_{s}b_{V_{0}}(V_{u})\,\mathrm{d}u\,\mathrm{d}s}{\mathcal{G}_{T}}+\frac{\int_{0}^{T}D_{s}\mathcal{G}_{T}\left(\int_{0}^{T}b_{V_{0}}(V_{u})\,\mathrm{d}u\right)\,\mathrm{d}s}{\mathcal{G}_{T}^{2}}\right)\right]$
$+\mathrm{e}^{-rT}\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\left(\int_{0}^{T}a_{V_{0}}(V_{s})\,\mathrm{d}W_{s}\left(\frac{W_{T}}{\mathcal{G}_{T}}-\frac{\int_{0}^{T}D_{s}\mathcal{G}_{T}}{\mathcal{G}_{T}^{2}}\right)-\frac{\int_{0}^{T}a_{V_{0}}(V_{t})+\int_{0}^{T}D_{t}a_{V_{0}}(V_{s})\,\mathrm{d}W_{s}\,\mathrm{d}t}{\mathcal{G}_{T}}\right)\right],$
$\mathcal{H}=$ $\mathrm{e}^{-rT}\mathbb{E}_{\mathbb{Q}}\left[f(S_{t})\left(\frac{W_{T}\int_{0}^{T}b_{H}(V_{s})\,\mathrm{d}s}{\mathcal{G}_{T}}-\frac{\int_{0}^{T}\int_{0}^{T}D_{s}b_{H}(V_{u})\,\mathrm{d}u\,\mathrm{d}s}{\mathcal{G}_{T}}+\frac{\int_{0}^{T}D_{s}\mathcal{G}_{T}\left(\int_{0}^{T}b_{H}(V_{u})\,\mathrm{d}u\right)\,\mathrm{d}s}{\mathcal{G}_{T}^{2}}\right)\right]$
$+\mathrm{e}^{-rT}\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\left(\int_{0}^{T}a_{H}(V_{s})\,\mathrm{d}W_{s}\left(\frac{W_{T}}{\mathcal{G}_{T}}-\frac{\int_{0}^{T}D_{s}\mathcal{G}_{T}}{\mathcal{G}_{T}^{2}}\right)-\frac{\int_{0}^{T}a_{H}(V_{t})+\int_{0}^{T}D_{t}a_{H}(V_{s})\,\mathrm{d}W_{s}\,\mathrm{d}t}{\mathcal{G}_{T}}\right)\right],$
where
$D_{t}a_{\theta}(V_{s})$ $=D_{t}\partial_{\theta}(V_{s})$
$=\partial_{\theta}\sigma^{\prime}(V_{s})D_{t}V_{s}$
$=\sigma^{\prime\prime}(V_{s})D_{t}V_{s}\partial_{\theta}V_{s}+\sigma^{\prime}(V_{s})\partial_{\theta}D_{t}V_{s}$
and
$D_{t}b_{\theta}(V_{s})$ $=D_{t}\partial_{\theta}\left(r-\frac{1}{2}\sigma^{2}(V_{s})\right)$
$=\partial_{\theta}D_{t}\frac{-\sigma^{2}(V_{s})}{2}$
$=-\partial_{\theta}\sigma(V_{s})\sigma^{\prime}(V_{s})D_{t}V_{s}$
$=-\left(\sigma^{\prime}(V_{s})^{2}+\sigma(V_{s})\sigma^{\prime\prime}(V_{s})\right)D_{t}V_{s}\partial_{\theta}V_{s}-\sigma(V_{s})\sigma^{\prime}(V_{s})\partial_{\theta}(D_{t}V_{s}).$
@@ 4.1 $\alpha$RFSV model
To calculate the considered Greeks, we need to calculate $D_{s}V_{r}$, $D_{t}D_{s}V_{r}$, $G(t,T)$, $D_{s}G(t,T)$, $D_{t}D_{s}G(t,T)$ and the derivatives of $V_{t}$ and $DV_{t}$ with respect to $V_{0}$ and $H$.
@@
$D_{t}D_{s}V_{r}=\rho^{2}\xi^{2}K_{H}(r,s)K_{H}(r,t)V_{r}\mathbf{1}_{[0,r]}(s\vee t).$
Moreover, from (37) we have that
$a_{V_{0}}(V_{t})=\partial_{V_{0}}V_{t}=\frac{V_{t}}{V_{0}},\quad b_{V_{0}}(V_{s})=\partial_{V_{0}}\left(r-\frac{V_{s}^{2}}{2}\right)=\frac{-V_{s}^{2}}{V_{0}}.$
Since $\sigma(x)=x$, the formulas needed to compute $\mathcal{V}$ are quite simple. Indeed,
$D_{t}a_{V_{0}}(V_{s})=\partial_{V_{0}}D_{t}V_{s}=\frac{1}{V_{0}}\rho\xi V_{s}K_{H}(s,t)\mathbf{1}_{[0,s]}(t).$
Hence,
$\int_{0}^{T}a_{V_{0}}(V_{t})+\int_{0}^{T}D_{t}a_{V_{0}}(V_{s})\,\mathrm{d}W_{s}\,\mathrm{d}t=\frac{1}{V_{0}}\left(\int_{0}^{T}V_{t}\,\mathrm{d}t+\int_{0}^{T}V_{s}\kappa_{s}\,\mathrm{d}W_{s}\right).$
In the same way,
$D_{t}b_{V_{0}}(V_{s})$ $=$ $-D_{t}V_{s}\cdot\frac{V_{s}}{V_{0}}-V_{s}\partial_{V_{0}}D_{t}V_{s}$
$=$ $-\frac{2}{V_{0}}D_{t}V_{s}$
$=$ $\frac{2}{V_{0}}\rho\xi V_{s}K_{H}(s,t)\mathbf{1}_{[0,s]}(t).$
Hence,
$\int_{0}^{T}b_{V_{0}}(V_{u})\,\mathrm{d}u=-\frac{1}{V_{0}}\int_{0}^{T}V_{u}^{2}\,\mathrm{d}u,\quad\int_{0}^{T}\int_{0}^{T}D_{s}b_{V_{0}}(V_{u})\,\mathrm{d}u\,\mathrm{d}s=-\frac{2}{V_{0}}\rho\xi\int_{0}^{T}V_{u}\kappa_{u}\,\mathrm{d}u.$
Regarding the derivatives with respect to $H$ we have that
$a_{H}(V_{t})=\partial_{H}V_{t}=V_{t}\cdot\left(\xi Y^{\prime}_{t}-\alpha\xi^{2}t^{2H}\ln(t)\right)$
where
$Y^{\prime}_{t}=\int_{0}^{t}\partial_{H}K_{H}(t,s)\,\mathrm{d}W_{s}.$
Hence,
$D_{s}a_{H}(V_{u})=\partial_{H}D_{s}V_{u}=$ $\rho\xi\partial_{H}V_{r}K_{H}(u,s)\mathbf{1}_{[0,u]}(s)+\rho\xi V_{u}\partial_{H}K_{H}(u,s)\mathbf{1}_{[0,u]}(s)$
$=$ $\left[\rho\xi V_{u}\left(\xi Y_{u}^{\prime}-\alpha\xi^{2}u^{2H}\ln(u)\right)K_{H}(u,s)+\rho\xi V_{u}\partial_{H}K_{H}(u,s)\right]\mathbf{1}_{[0,u]}(s).$
This implies that
$\int_{0}^{T}a_{H}(V_{t})+\int_{0}^{T}D_{t}a_{H}(V_{s})\,\mathrm{d}W_{s}\,\mathrm{d}t$
$=$ $\int_{0}^{T}V_{t}\cdot\left(\xi Y^{\prime}_{t}-\alpha\xi^{2}t^{2H}\ln(t)\right)\,\mathrm{d}t+\rho\xi\int_{0}^{T}\left[V_{s}(\xi Y_{s}^{\prime}-\alpha\xi^{2}s^{2H}\ln(s))\kappa_{s}+V_{s}\kappa^{\prime}_{s}\right]\,\mathrm{d}W_{s}.$
Similarly, for $b_{H}$ we have
$b_{H}(V_{t})=-\partial_{H}\frac{V_{t}^{2}}{2}=-V_{t}\partial_{H}V_{t}=-V_{t}^{2}(\xi Y_{t}^{\prime}-\alpha\xi^{2}t^{2H}\ln(t))$
and
$\int_{0}^{T}\int_{0}^{T}D_{s}b_{H}(V_{u})\,\mathrm{d}u\,\mathrm{d}s=-$ $\int_{0}^{T}\int_{0}^{T}\partial_{H}V_{u}\cdot D_{s}V_{u}-V_{u}\cdot\partial_{H}(D_{s}V_{u})\,\mathrm{d}u\,\mathrm{d}s$
=-= - $\rho\xi\int_{0}^{T}V_{u}^{2}\left[2(\xi Y_{u}^{\prime}-\alpha\xi^{2}u^{2H}\ln(u))\kappa_{u}+\kappa^{\prime}_{u}\right]\,\mathrm{d}u$
@@
Using these formulas we have
@@
$=\int_{0}^{T}V_{t}\,\mathrm{d}t+\rho\xi\left(\int_{0}^{T}V_{r}\kappa_{r}\,\mathrm{d}W_{r}-\int_{0}^{T}V_{r}^{2}\kappa_{r}\,\mathrm{d}r\right).$ (38)
Differentiating this last expression, we obtain
@@
$=2\rho\xi\int_{0}^{T}V_{t}\kappa_{t}\,\mathrm{d}t+\rho^{2}\xi^{2}\left(\int_{0}^{T}V_{r}\kappa_{r}^{2}\,\mathrm{d}W_{r}-2\int_{0}^{T}V_{r}^{2}\kappa_{r}^{2}\,\mathrm{d}r\right)$ (39)
and finally,
$\int_{0}^{T}\int_{0}^{T}\int_{0}^{T}D_{t}D_{s}G(u,t)\,\mathrm{d}u\,\mathrm{d}s\,\mathrm{d}t=$ $2\rho^{2}\xi^{2}\int_{0}^{T}\kappa_{u}^{2}V_{u}\,\mathrm{d}u$
$+$ $\rho^{2}\xi^{2}\int_{0}^{T}\kappa_{u}^{2}V_{u}\,\mathrm{d}u+\rho^{3}\xi^{3}\int_{0}^{T}\kappa_{u}^{3}V_{u}\,\mathrm{d}W_{u}$
-- $4\rho^{3}\xi^{3}\int_{0}^{T}\kappa_{u}^{3}V_{u}^{2}\,\mathrm{d}u$
$=$ $3\rho^{2}\xi^{2}\int_{0}^{T}\kappa_{u}^{2}V_{u}\,\mathrm{d}u-\rho^{3}\xi^{3}\left(\int_{0}^{T}\kappa_{u}^{3}V_{u}\,\mathrm{d}W_{u}-4\int_{0}^{T}\kappa_{u}^{3}V_{u}^{2}\right).$
In conclusion, the tools needed to compute the Greeks for the $\alpha$RFSV model are the following:
  • $\mathcal{G}_{T}=\int_{0}^{T}V_{t}\,\mathrm{d}t+\rho\xi\left(\int_{0}^{T}V_{r}\kappa_{r}\,\mathrm{d}W_{r}-\int_{0}^{T}V_{r}^{2}\kappa_{r}\,\mathrm{d}r\right)$,
  • $\int_{0}^{T}D_{s}\mathcal{G}_{T}\,\mathrm{d}s=2\rho\xi\int_{0}^{T}V_{t}\kappa_{t}\,\mathrm{d}t+\rho^{2}\xi^{2}\left(\int_{0}^{T}V_{r}\kappa_{r}^{2}\,\mathrm{d}W_{r}-2\int_{0}^{T}V_{r}^{2}\kappa_{r}^{2}\,\mathrm{d}r\right)$,
  • $\int_{0}^{T}\int_{0}^{T}D_{t}D_{s}\mathcal{G}_{T}\,\mathrm{d}s\,\mathrm{d}t=3\rho^{2}\xi^{2}\int_{0}^{T}\kappa_{u}^{2}V_{u}\,\mathrm{d}u-\rho^{3}\xi^{3}\left(\int_{0}^{T}\kappa_{u}^{3}V_{u}\,\mathrm{d}W_{u}-4\int_{0}^{T}\kappa_{u}^{3}V_{u}^{2}\right)$,
  • $a_{V_{0}}(V_{t})=\frac{V_{t}}{V_{0}}$,
  • $b_{V_{0}}(V_{t})=-\frac{V_{t}^{2}}{2}$,
  • $\int_{0}^{T}a_{V_{0}}(V_{t})+\int_{0}^{T}D_{t}a_{V_{0}}(V_{s})\,\mathrm{d}W_{s}\,\mathrm{d}t=\frac{1}{V_{0}}\left(\int_{0}^{T}V_{t}\,\mathrm{d}t+\int_{0}^{T}V_{s}\kappa_{s}\,\mathrm{d}W_{s}\right)$,
  • $\int_{0}^{T}\int_{0}^{T}D_{s}b_{V_{0}}(V_{u})\,\mathrm{d}u\,\mathrm{d}s=\frac{-2}{V_{0}}\rho\xi\int_{0}^{T}V_{u}\kappa_{u}\,\mathrm{d}u$,
  • $a_{H}(V_{t})=V_{t}\cdot\left(\xi Y^{\prime}_{t}-\alpha\xi^{2}t^{2H}\ln(t)\right)$, where $Y^{\prime}_{t}=\int_{0}^{t}\partial_{H}K_{H}(t,s)\,\mathrm{d}W_{s}$,
  • $b_{H}(V_{t})=-V_{t}^{2}\cdot\left(\xi Y^{\prime}_{t}-\alpha\xi^{2}t^{2H}\ln(t)\right)$,
  • $\int_{0}^{T}a_{H}(V_{t})+\int_{0}^{T}D_{t}a_{H}(V_{s})\,\mathrm{d}W_{s}\,\mathrm{d}t$
    $=\int_{0}^{T}V_{t}\cdot\left(\xi Y^{\prime}_{t}-\alpha\xi^{2}t^{2H}\ln(t)\right)\,\mathrm{d}t+\rho\xi\int_{0}^{T}\left[V_{s}(\xi Y_{s}^{\prime}-\alpha\xi^{2}s^{2H}\ln(s))\kappa_{s}+V_{s}\kappa^{\prime}_{s}\right]\,\mathrm{d}W_{s},$
  • $\int_{0}^{T}\int_{0}^{T}D_{s}b_{H}(V_{u})\,\mathrm{d}u\,\mathrm{d}s=-\rho\xi\int_{0}^{T}V_{u}^{2}\left[2(\xi Y_{u}^{\prime}-\alpha\xi^{2}u^{2H}\ln(u))\kappa_{u}+\kappa^{\prime}_{u}\right]\,\mathrm{d}u$.
These are all ingredients necessary to plug into the Greeks formulas in Proposition 4.1.
@@
In Figure 1 we can see a convergence of the Delta for the $\alpha$RFSV model with $H=0.15$, $\alpha=1$ (rBergomi), $V_{0}=0.354$, $\xi=0.212$ and $\rho=-0.756$. Numerically, the kernel (23) is considered with $\varepsilon=10^{-6}$. Market values are $S_{0}=100$ (dollars) and $r=0.05$, option parameters are strike price $K=100$ (dollars) and maturity $T=1$ (year). Monte-Carlo (MC) simulation point estimate together with the 99% confidence interval (CI) is plotted as a dependence on the number of simulations (NS). In particular $\Delta_{\text{call}}=0.3399\pm 0.0051$ and $\Delta_{\text{put}}=-0.9220\pm 0.0156$. Although we have no analytical $\alpha$RFSV formula for Delta, we plot the horizontal dashed line which is the numerical mean at the end (for the largest number of simulations).
Figure 1: Convergence of the $\alpha$RFSV Delta in the case $H=0.15$.
In Figure 2 we depict the dependence of the $\alpha$RFSV Delta on $H$, both for the call and put option case, with remaining parameters left unchanged.
Figure 2: Dependence of the $\alpha$RFSV Delta on the Hurst parameter $H$.